-2x+11+6x
-5(1-5x)+5(-8x-2)= -4-8x
-2n-(9-10n)
-3x-9+15x
-5n+3(6+7n)

Answer by checkley79(3341)

-2x+11+6x
4x+11 ans.
------------------
-5(1-5x)+5(-8x-2)= -4-8x
-5+25x-40x-10=-4-8x
-15x-15=-4-8x
-15x+8x=-4+15
-7x=11
x=11/-7 ans.
--------------------------
-2n-(9-10n)
-2n-9+10n
8n-9 ans.
---------------------------
I'll leave the last 2 for you to practice with.
-3x-9+15x
-5n+3(6+7n)